OpenStudy (anonymous):

Find dy/dx of y=x(lnx)^(sin^2x) using log differentiation.

OpenStudy (anonymous):

lny = ln(x(lnx)^(sin^2x)) Can someone help me distribute the ln on the right side I am having trouble

OpenStudy (anonymous):

lny = ln x + (sin^2 x) ln (lnx) y`/y = 1/x + sin2x (ln(lnx)) + sin^2x/(x lnx)

OpenStudy (anonymous):

ln x + (sin^2 x) ln (lnx) can you try to explain this step more?